You are employed for a consulting company as a con
Answer #1

pE = pE0 - 1/n log Q

E = E0 - 0.0592 / n log Q

For redcution of nitrate

Q = pN2 / [NO3-]^2 [ [H+]12

Given pH = 8 therefore [H+] = 10^-8

Q = 1/ (4 X 10^-4 )^2 (10^-8)^12

E0 = 1.25

E = 1.25 - 0.0592 / 10 log(1/ (4 X 10^-4 )^2 (10^-8)^12)

E =1.869 V

pE = E / 2.303RT / F

F= 96485

R = 8.314

T = 298

pE = 0.1.869 / 0.0592 = 0.619 / 0.0592 = 31.57

b) pE for reduction of oxygen as a function of the partial pressure of oxygen

E = E0 -0.0592 / n log Q

pE = pE0 - 1/n log Q

Q = 1 / [H+]^4 X partial pressure of oxygen